In a right triangle, one side of teh right angle is located on the X-axis; teh other side of the right angle is located on teh y-ais. The hypotenuse of this triangle is formed by a segment of one line passing through the point M(2,3). The area of this right triangle is 12 square units.
What is the equation of the line passing through the point M?

Answer by checkley71(8403)

the other two points are (4,0) & (0,6) which creates a triangle with a 4 unit base and a height of 6 thus the area id 4/2*6=12 or 2*6=12.
and the equation of the line pasisng through M(2,3) is Y=-3/2X+6